Can you please show the work for this step by step? I would greatly appreciate it I keep getting off by a few numbers. Thank you!Derek will deposit $2,091.00 per year into an account starting today and ending in year 13.00. The account that earns 9.00%.
Answer #1

As per Annuity Formula

Future Value = (C/r)[(1+r)n-1]

C = annual Payment per year = $2091

r = interest rate annually =9%

n = number of years =13 years

Future Value = (C/r)[(1+r)n-1]

Since the amount deposited at time t =0

Therefore Future Value = (C/r)[(1+r)n-1] + C(1+r)n

=$(2091/0.09)[(1.09)13-1] + 2091(1.09)13

=$54406.12